OpenStudy (anonymous):

Naveena used a 15$ coupon for dinner with her mother. She paid 17.70 after the coupon was applied. Which equation represents this problem? Let c represent the original cost of the dinner.

OpenStudy (anonymous):

C= 15$ + 17.70

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

choices are 15c=17.70 or c-15=17.70 or c/15 =17.70 or c+15=17.70

OpenStudy (anonymous):

B I think..